OpenStudy (anonymous):

Find the power series representation for the function and determine the interval of convergence f(x)=2/(3-x)

OpenStudy (anonymous):

Find the nth derivative: \[D \left( \frac{2}{3-x}\right)=\frac{2}{(3-x)^2}; D^2 \left( \frac{2}{(3-x)}\right)=\frac{2(2)}{(3-x)^3} \] \[\implies D^n \left( \frac{2}{3-x}\right) \left( \frac{2n!}{(3-x)^{n+1}}\right)\] Evaluating this at x=0 gives: \[\frac{2n!}{3^{n+1}}\] So: \[f(x)=\sum_{n=0}^{\infty} \frac{f^{(n)}(x)(x-a)^n}{n!}\] I'm assuming you want the McLaurin series where a=0 so: \[f(x)=\sum_{n=0}^{\infty} \frac{2x^n}{3^{n+1}}\] Then apply the ratio test: \[\lim_{k \rightarrow \infty}\left| \frac{a_{k+1}}{a_k} \right|<|1|\] This gives you the interval. Then plug in the end points for x to find if its less than or less than or equal to.